Circular No. 5848

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET MUELLER (1993p)

The following precise positions have been reported:

1993 UT α_{2000} δ_{2000} Observer
Aug. 18.93596 1^{h}44^{m}06.^{s}62 +49^{o}55′00′′.0 Tichý
18.94753 1 44 06.23 +49 55 01.9  _{”}
19.29623 1 43 53.63 +49 55 47.0 Tatum
19.29785 1 43 53.58 +49 55 47.2  _{”}
19.29925 1 43 53.52 +49 55 47.4  _{”}

M. Tichý (Kleť). 0.63-m Maksutov telescope. Communicated by J. Tichá. J. B. Tatum and D. D. Balam (Climenhaga Observatory). 0.5-m reflector

+ CCD. Comet well condensed, coma diameter 7′′; fanshaped tail 30′′ long in p.a. 270^{o}.

Preliminary parabolic orbital elements by the undersigned:

T = 1994 Apr. 5.04 TT  ω = 261.^{o}27
Ω = 193.95 }2000.0
q = 0.9835 AU  i = 105.58
1993 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Aug. 11 1^{h}47.^{m}74 +49^{o}33.^{′}6 3.290 3.493 93.^{o}0 16.^{o}8 14.0
16 1 45.71 +49 48.0 3.163 3.438  97.1 17.0 13.9
21 1 42.79 +49 59.2 3.036 3.382 101.2 17.1 13.7
26 1 38.90 +50 06.4 2.911 3.327 105.5 17.0 13.5
31 1 33.98 +50 08.3 2.787 3.270 109.8 16.9 13.4
Sept. 5 1 27.96 +50 03.5 2.666 3.214 114.2 16.6 13.2
10 1 20.82 +49 50.5 2.548 3.157 118.7 16.3 13.0
15 1 12.55 +49 27.2 2.434 3.100 123.1 15.8 12.8
20 1 03.22 +48 51.5 2.325 3.043 127.5 15.2 12.7

SUPERNOVA 1993W IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of a

supernova on films obtained by E. W. Elst and D. Albanese with the 0.9-m Schmidt telescope on Aug. 18.06 and 19.05 UT. The supernova, of V = 17, is located at α = 23^{h}27^{m}05^{s}.78, δ = −3^{o}18′58′′.4 (equinox 1950.0), 1′′.8 east and 8′′.7 south of the center of an anonymous spiral galaxy of V = 16. A nearby star (V = 15.5) has end figures 01^{s}.86, 20′13′′.5.

1993 August 19 Brian G. Marsden